OpenStudy (anonymous):

slope intercept for this x=3 y=3/5

OpenStudy (anonymous):

Is that all the question said?

OpenStudy (anonymous):

find the slope intercept form for the line satifying the following condition x=3 and y=3/5 the slop intercept form the for the line is ??

OpenStudy (mayankdevnani):

x+y=18/5

OpenStudy (anonymous):

how did you get that pleasse explain

OpenStudy (mayankdevnani):

add x=3 and y=3/5 what you get

OpenStudy (anonymous):

There are many answers. The easiest one is: x=3, so x/3=1 y=3/5, so 5/3y=1 Then, x/3 = 5/3y. And so: x = 5y y = x/5

OpenStudy (mayankdevnani):

can you solve 3+3/5=?

OpenStudy (mayankdevnani):

@muntoo don't give direct answers

OpenStudy (mayankdevnani):

plz teach

OpenStudy (anonymous):

6/5=1.2

OpenStudy (mayankdevnani):

no

OpenStudy (anonymous):

3 + (3/5)

OpenStudy (mayankdevnani):